Mathematics 133 – Intensive Calculus for Science 1 Quiz 3 – September 23, 2005

Your Name:

Do all work on this sheet (front and back if needed). You have 20 minutes. No calculators on this quiz.

A) This is part of the graph of a periodic function of x. Give a possible formula for the function:

B) The following graph shows the graph of a polynomial function y = f(x) (the plot shows a big enough x-range so that we are seeing all relevant features). What is the smallest possible degree of the polynomial f(x)?

C) Which of the following two graphs shows y = \frac{x^{2}−9}{x^{2}+1}. Explain how you can tell.